# Constraining the Stellar-to-Halo Mass Relation with Galaxy Clustering and Weak Lensing from DES Year 3 Data

### v2

@@ 7.1.1 SHMR
We compare our results with independent constraints from the literature, as well as with the prediction of Behroozi et al. (2013), Moster et al. (2013), Kravtsov et al. (2018), and Thornton et al. (2024) in Figure 9. The filled and open blue triangles present the results from CFHTLenS (Velander et al., 2013), where the samples they considered were flux-limited and were split into blue and red galaxies, respectively. The filled and open red triangles in the same plot present the constraints from SDSS (Mandelbaum et al., 2006) and galaxies are, respectively, split into early and late types. The magenta squares denote the average halo masses inferred from individual galaxy stellar mass samples using the weak lensing measurements in GAMA (Han et al., 2015) where the central galaxies of groups and clusters were almost volume-limited. We finally also compare our results with those from Zu & Mandelbaum (2015), where a HOD model similar to this work is used to jointly fit galaxy-galaxy lensing and galaxy clustering in SDSS. We note that all those studies were carried out in similar redshift ranges where the SHMR is only weakly dependent on redshift. In addition, we find that we do not have the constraining power to independently fit to our redshift bins, whereas by combining all data points we can obtain meaningful constraints on the SHMR model parameters. Overall, these constraints are broadly in agreement. We also note that all of these analyses make different assumptions in their models, priors and covariances, so it is not straightforward to compare the datasets directly beyond the high-level statement above.
@@ 7.2 Sensitivity to analysis choices
We examine three analysis choices in our pipeline and study the sensitivity of our main results to modifying them. These three variations are: 1) exclude galaxy clustering from the fit and only use galaxy-galaxy lensing, similar to Zacharegkas et al. (2022); 2) assume a different set of fiducial cosmological parameters; 3) use a different SHMR functional form Equation 2.1 in our modeling. All three tests are summarized in Figures 12 and 13.
@@ 7.2.1 Fitting $\gamma_{t}$ only
The model in this paper is developed to fit simultaneously $w$ and $\gamma_{t}$. When we exclude galaxy clustering from our model fitting, the resulting halo mass constraints do not change significantly for the lowest mass bins but become underestimated for the medium-to-high mass ranges, as shown in Figure 12 as the black points. This is related to the fact that galaxy clustering is significantly more sensitive to the satellite terms in the model than galaxy-galaxy lensing. Our results imply that with a complex galaxy-halo connection model like the one we have utilized in this paper, which has a lot of flexibility, careful choices of parameter priors are necessary. For instance, in the absence of $w$, as we have also seen in Zacharegkas et al. (2022), the central and satellite 2-halo terms in $\gamma_{t}$ can be indistinguishable by the fitting algorithm and priors might need to be applied to the satellite fraction to break this behavior. When clustering is included, this is mostly avoided without additional priors because of how sensitive $w$ is to the satellite galaxies.
@@ 7.2.2 Assumption of fiducial cosmology
As described in Section 1, throughout the paper we have assumed a hybrid cosmology of DES Y3 and Planck 2018 best-fit. We now test whether assuming only the Planck 2018 best-fit cosmology (Planck Collaboration, 2020) changes the main result of the paper. The resulting changes in mean halo mass for each bin as well as the SHMR constraints are shown in Figures 12 and 13, respectively. We find that the change in cosmology has negligible effect on both.
@@ 8 Summary and discussion
In order to analyze our stellar-mass sample, we developed a HOD model that is an update from that used in Zacharegkas et al. (2022). The new model accounts for the stellar mass information of the galaxies and further incorporate galaxy clustering into the framework. The final goal of this work is to fit the data across all redshift and stellar mass range in order to constrain the stellar-to-halo mass relation for this sample. Our main findings are the following:
@@
  • In order to obtain good data fits to all scales, especially in low-stellar mass bins, we needed to update how we model the 1-to-2-halo transition scales compared to what was utilized in Zacharegkas et al. (2022). Our updated model required a more careful treatment of those intermediate scales, which implies that a much better understanding of the galaxy-halo connection is required in order to describe our high-precision measurements.
@@
Although our model is flexible enough to fit the data, we emphasize here that it can be further improved. The 1-to-2 halo regime, as mentioned above, had to be modified specifically to fit the low mass bins and we expect that further modification, e.g. by adding halo exclusion, could further add freedom to the model to fit the data better. In addition, galaxy clustering is very sensitive to the modeling choices, especially those concerning the satellite terms. Decisions regarding the halo mass function and matter power spectrum on the nonlinear and semi-linear regime can also make a difference in the data fits. In short, we find that practically each model component and every model complexity ultimately leads to a framework that needs to be calibrated specifically to fit the data at hand in the HOD framework we presented for such high-precision data. This can be seen by how swapping the SHMR, for instance, in our test in Section 7.2.3, slightly changes our final results. In order to improve the fits in that case we would need to recalibrate our model, given the new model parameters and reduced flexibility from the new SHMR. In addition, we discuss how in Section 7.2.1 our results are also changed when only lensing is considered. When adding new observables we see significant changes to the fits and the model also needs to be recalibrated. We saw a similar, albeit simpler, manifestation of this in Zacharegkas et al. (2022) in the case of the redMaGiC sample where our constraints immediately improved by adding the galaxy number density as part of the data vector.
@@
The fact that empirical models such as the HOD are being gradually made more complex as the data sets become more constraining could be a manifestation of the need for a more complete and physics-oriented approach to studying the galaxy-halo connection. This paper is an example of this, as we had to improve upon our previous work (Zacharegkas et al., 2022) significantly in order to be able to describe our new dataset. Today, alongside the empirical models, such as the HOD, Conditional Luminosity Function (CLF) or abundance matching, techniques such as forward-modeling are commonly utilized (e.g., Hearin et al., 2017; Chittenden & Tojeiro, 2022; Hahn et al., 2023). The latter are based on producing synthetic galaxy catalogs to make observations on, which are then compared to real data. The simulations that produce the synthetic galaxy populations can incorporate the physics of interest, albeit limitations exist such as resolution and simulation volume, and they can be a lot more computationally expensive, and hard to fully interpret, than empirical models that depend on a few parameters and are based on a few transparent assumptions.
@@ Appendix A A generalized HOD model given joint redshift-stellar mass distribution
In this appendix we present a more general case of our fiducial model, where the full 2D distribution of redshift and stellar mass for each lens bin is considered, similar to Leauthaud et al. (2011); Zu & Mandelbaum (2015). Our fiducial model is a special case in which 1) the redshift and stellar mass distributions are uncorrelated and 2) the stellar mass distribution is approximated as a tophat function.
@@
[…]
given a stellar mass range $[M_{\star,1},M_{\star,2}]$ and a redshift range $[z_{\ell,1},z_{\ell,2}]$. Henceforth, we will be denoting joint probability distributions by the letter ”$p$”, and 1D distributions by ”$n$”.
@@
×$\qquad\times\int dM_{\star}\;p(M_{h},M_{\star})n_{\star}(\log M_{\star}|z_{\ell})\;,$ (49)
where $p(M_{h},M_{\star})$ is the probability of finding a galaxy with $M_{\star}$ in a halo of mass $M_{h}$, $n_{\star}(M_{\star}|z_{\ell})$ is the distribution of stellar mass at redshift $z_{\ell}$, and where we have defined $n_{\star}(\log M_{\star})=[M_{\star}/\log(e)]n_{\star}(M_{\star})$. Also, in the above, $dn/dM_{h}$ is the Halo Mass Function (HMF) for which we use Tinker et al. (2008) in the paper, and $n_{g}$ is the average number density of galaxies. In addition, in Equation (A) above we can write
@@
[…]
×$\quad\times\langle N(M_{\star}|M_{h})\rangle n_{\star}(\log M_{\star}|z_{\ell})$
[…]
×$\quad\times\langle N(M_{\star}|M_{h})\rangle\frac{p(\log M_{\star},z_{\ell})}{n_{\ell}(z_{\ell})}\;,$ (51)
@@
×$\quad\times\exp\left[-\left(\frac{\log M_{\star}-\log f_{\rm SHMR}(M_{h})}{\sigma_{\log M_{\star}}}\right)^{2}\right]\;.$ (52)
@@ Appendix B Validating the stellar mass catalog
We run two validation tests for our GalPro-generated stellar mass catalog based on Gneiting et al. (2007), which we summarize below. In order to perform these tests and create the plots we present in this section, we performs a $80/20$ split to create the training and validation sets, respectively, from the matched DES Y3 Deep Field and COSMOS2020 catalogs (see Section 4.1.1). GalPro is trained with the training set and the tests below are done with the validation set. See also Mucesh et al. (2020) for more details regarding the validation tests. More specifically, we test for probabilistic calibration, using the Probability Integral Transform (PIT), and for marginal calibration, by comparing the average predictive CDF and the true empirical CDF. We present the results, while briefly describing the process, for both tests below.
@@
    The results of the above tests on our model are shown in Figure 15. The PIT distributions follow closely the uniform. In addition, three statistics listed in the caption of the plot are given to quantitatively assess uniformity: the Kullback–Leibler (Kullback & Leibler, 1951, KL) divergence, the Kolmogorov–Smirnov (Kolmogorov, 1992, KS) test, and the Cramér-von Mises (Cramér, 1928, CvM) test. For each of these tests, values close to zero would indicate that our PIT distribution closely matches the Uniform distribution. We conclude that our model passes this test.
@@
    Although the PIT test is a necessary condition for the marginal PDF to be valid, it is usually not enough to make concrete statements about how well a PDF is calibrated. In fact, it has been shown (Hamill, 2001) that biased distributions can lead to uniformly distributed PIT’s. This test is thus not sufficient and we also perform the additional test below.
@@
    where we have defined the predictive CDF per galaxy as $F_{i}$ for all $n$ test galaxies $i=1,2,...,n$, and where
@@
[…]
@@ Appendix C Computing the LSS weights
To prevent the network from overfitting, it is constructed with k-fold cross-validation, which works in the following way: The NSIDE = $4096$ maps are re-binned into a coarser grid of NSIDE = $16$ (with a resolution of about 4 degrees). We then randomly divide these cells into $k$ equal-area groups. To derive the weights for a given fold $k$, we train the neural network on the other folds, using fold $k$ as a validation sample (the training halts when the training metric no longer improves on the validation set). This cross-validation scheme will only work to prevent overfitting on scales below the resolution defined by NSIDE split, in this case around 4 degrees. As also mentioned in Sánchez et al. (2022), testing the corrected and uncorrected galaxy clustering on log-normal mock catalogs demonstrated no overfitting from this method at scales below 1 degree, and an impact of around 5% overfitting at scales of 2 degrees. We thus keep the galaxy clustering analysis in this work to angular scales below 1 degree to be conservative.
@@ Appendix D Validation of covariance matrix
As discussed in Section 5.2, our main analysis uses a Jackknife covariance, but we also calculate a theoretical covariance matrix via the CosmoCov (Krause & Eifler, 2017; Fang et al., 2020) package. Figure 17 shows the comparison between the Jackknife and CosmoCov error bars (the square root of the diagonal of the covariance matrix) in the case of lens-source bin combination $(\ell,m,s)=(2,1,4)$, for $\gamma_{t}$ (upper panel) and $w$ (lower panel); we use that bin for demonstration here but we see similar trends in all bins. In the left plot we also show the line $\propto 1/\theta$ which shows the expected behavior of the covariance if it were shape noise dominated. As we see in that plot, $\gamma_{t}$ is dominated by shape noise on small scales. This also implies that the cross-covariance between $\gamma_{t}$ and $w$ in a single bin and across different redshift and stellar-mass bins is weak. Therefore, we choose to ignore the cross-covariance in this work.
@@
The agreement between Jackknife and theory is not perfect. However, this is not very surprising given the differences in the HOD theory used by CosmoCov compared to this study, as the one implemented in the former is based on Zheng et al. (2005); Zehavi et al. (2011). Nevertheless, the differences we see in the covariance are not expected to have a significant effect on our data fits and we proceed to use the Jackknife covariace matrix in Section 6 while performing the likelihood analysis.

### v3

@@ 7.1.1 SHMR
We compare our results with independent constraints from the literature, as well as with the prediction of Behroozi2013, Moster2013, Kravtsov2018, and Thornton2023 in Figure 9. The filled and open blue triangles present the results from CFHTLenS (velander2013), where the samples they considered were flux-limited and were split into blue and red galaxies, respectively. The filled and open red triangles in the same plot present the constraints from SDSS (Mandelbaum2006) and galaxies are, respectively, split into early and late types. The magenta squares denote the average halo masses inferred from individual galaxy stellar mass samples using the weak lensing measurements in GAMA (Han2015) where the central galaxies of groups and clusters were almost volume-limited. We finally also compare our results with those from zu2015, where a HOD model similar to this work is used to jointly fit galaxy-galaxy lensing and galaxy clustering in SDSS. We note that all those studies were carried out in similar redshift ranges where the SHMR is only weakly dependent on redshift. In addition, we find that we do not have the constraining power to independently fit to our redshift bins, whereas by combining all data points we can obtain meaningful constraints on the SHMR model parameters. Overall, these constraints are broadly in agreement. We emphasize, however, that all of these analyses make different assumptions in their models, priors and covariances, use different cosmological models, halo mass definitions, and they even differ in how they measured and reported stellar masses; thus, it is not straightforward to compare the datasets directly to make statements beyond the high-level comparisons we comment on here. To correct and account for all these differences in the reported results would be complicated and outside the scope of this work. We have, however, made sure that we convert all halo masses to the same mass definition, in particular $M_{\rm 200c}$. Below we briefly comment further on these comparisons.
Our model constraints are in excellent agreement with Kravtsov2018, whose formulation of the SHMR is very similar to this work. In that paper, the authors used abundance matching and re-analyzed data of stellar masses of clusters in SDSS to constrain the stellar mass-halo mass relation. They argue that previous studies that reported similar constraints using abundance matching underestimated the stellar mass in clusters. That could explain why compared to Behroozi2013 and Moster2013 we find overall agreement with our constraints on the small masses, while there is tension at higher masses $\log(M_{\rm h}^{\rm cen}/M_{\odot})\gtrsim 12$. This is an interesting observation, further investigation of which we leave for future work. Our results are also in good agreement with zu2015 who used a similar HOD framework to derive their constraints with galaxy-galaxy lensing and galaxy clustering. On the small-mass regime we find good agreement with Thornton2023 within 1-$\sigma$, where the SHMR was constrained using galaxy-galaxy lensing around photometrically selected dwarf galaxies. Our constraints demonstrate a different shape for the SHMR compared to velander2013, who utilized weak lensing to constrain the galaxy-halo connection. A key difference between that study and this work is that our model is more flexible and we have combined lensing with galaxy clustering which adds significantly more constraining power. In Mandelbaum2006 the authors used weak gravitational lensing to infer the galaxy-dark matter relation; our results are broadly in agreement within 1-$\sigma$, however we note that in this work we have made various different assumptions, including the assumed cosmological parameter values.
@@ 7.2 Sensitivity to analysis choices
We examine three analysis choices in our pipeline and study the sensitivity of our main results to modifying them. These three variations are: 1) exclude galaxy clustering from the fit and only use galaxy-galaxy lensing, similar to Zacharegkas2022; 2) assume a different set of fiducial cosmological parameters; 3) use a different SHMR functional form Equation 2.1 in our modeling. All three tests are summarized in Figures 12 and 13.
@@ 7.2.1 Fitting $\gamma_{t}$ only
The model in this paper is developed to fit simultaneously $w$ and $\gamma_{t}$. When we exclude galaxy clustering from our model fitting, the resulting halo mass constraints do not change significantly for the lowest mass bins but become underestimated for the medium-to-high mass ranges, as shown in Figure 12 as the black points. This is related to the fact that galaxy clustering is significantly more sensitive to the satellite terms in the model than galaxy-galaxy lensing. Our results imply that with a complex galaxy-halo connection model like the one we have utilized in this paper, which has a lot of flexibility, careful choices of parameter priors are necessary. For instance, in the absence of $w$, as we have also seen in Zacharegkas2022, the central and satellite 2-halo terms in $\gamma_{t}$ can be indistinguishable by the fitting algorithm and priors might need to be applied to the satellite fraction to break this behavior. When clustering is included, this is mostly avoided without additional priors because of how sensitive $w$ is to the satellite galaxies.
@@ 7.2.2 Assumption of fiducial cosmology
As described in Section 1, throughout the paper we have assumed a hybrid cosmology of DES Y3 and Planck 2018 best-fit. We now test whether assuming only the Planck 2018 best-fit cosmology (Planck2018VI) changes the main result of the paper. The resulting changes in mean halo mass for each bin as well as the SHMR constraints are shown in Figures 12 and 13, respectively. We find that the change in cosmology has negligible effect on both.
@@ 8 Summary and discussion
In order to analyze our stellar-mass sample, we developed a HOD model that is an update from that used in Zacharegkas2022. The new model accounts for the stellar mass information of the galaxies and further incorporate galaxy clustering into the framework. The final goal of this work is to fit the data across all redshift and stellar mass range in order to constrain the stellar-to-halo mass relation for this sample. Our main findings are the following:
@@
  • In order to obtain good data fits to all scales, especially in low-stellar mass bins, we needed to update how we model the 1-to-2-halo transition scales compared to what was utilized in Zacharegkas2022. Our updated model required a more careful treatment of those intermediate scales, which implies that a much better understanding of the galaxy-halo connection is required in order to describe our high-precision measurements.
@@
Although our model is flexible enough to fit the data, we emphasize here that it can be further improved. The 1-to-2 halo regime, as mentioned above, had to be modified specifically to fit the low mass bins and we expect that further modification, e.g. by adding halo exclusion, could further add freedom to the model to fit the data better. In addition, galaxy clustering is very sensitive to the modeling choices, especially those concerning the satellite terms. Decisions regarding the halo mass function and matter power spectrum on the nonlinear and semi-linear regime can also make a difference in the data fits. In short, we find that practically each model component and every model complexity ultimately leads to a framework that needs to be calibrated specifically to fit the data at hand in the HOD framework we presented for such high-precision data. This can be seen by how swapping the SHMR, for instance, in our test in Section 7.2.3, slightly changes our final results. In order to improve the fits in that case we would need to recalibrate our model, given the new model parameters and reduced flexibility from the new SHMR. In addition, we discuss how in Section 7.2.1 our results are also changed when only lensing is considered. When adding new observables we see significant changes to the fits and the model also needs to be recalibrated. We saw a similar, albeit simpler, manifestation of this in Zacharegkas2022 in the case of the redMaGiC sample where our constraints immediately improved by adding the galaxy number density as part of the data vector.
@@
The fact that empirical models such as the HOD are being gradually made more complex as the data sets become more constraining could be a manifestation of the need for a more complete and physics-oriented approach to studying the galaxy-halo connection. This paper is an example of this, as we had to improve upon our previous work (Zacharegkas2022) significantly in order to be able to describe our new dataset. Today, alongside the empirical models, such as the HOD, Conditional Luminosity Function (CLF) or abundance matching, techniques such as forward-modeling are commonly utilized (e.g., Hearin2017_halotools; Chittenden2022; ChangHoon2023). The latter are based on producing synthetic galaxy catalogs to make observations on, which are then compared to real data. The simulations that produce the synthetic galaxy populations can incorporate the physics of interest, albeit limitations exist such as resolution and simulation volume, and they can be a lot more computationally expensive, and hard to fully interpret, than empirical models that depend on a few parameters and are based on a few transparent assumptions.
@@ Appendix A A generalized HOD model given joint redshift-stellar mass distribution
In this appendix we present a more general case of our fiducial model, where the full 2D distribution of redshift and stellar mass for each lens bin is considered, similar to leauthaud2011; zu2015. Our fiducial model is a special case in which 1) the redshift and stellar mass distributions are uncorrelated and 2) the stellar mass distribution is approximated as a tophat function.
@@
$\int_{M_{\star,1}}^{M_{\star,2}}dM_{\star}\int_{z_{\ell,1}}^{z_{\ell,2}}dz_{\ell}\;p(M_{\star},z)=1,$ (47)
given a stellar mass range $M_{\star,1},M_{\star,2}]$ and a redshift range $z_{\ell,1},z_{\ell,2}]$. Henceforth, we will be denoting joint probability distributions by the letter ”pp”, and 1D distributions by ”nn”.
@@
$\qquad\times\int dM_{\star}\;p(M_{h},M_{\star})n_{\star}(\log M_{\star}|z_{\ell})\;,$ (49)
where $p(M_{h},M_{\star})$ is the probability of finding a galaxy with $M_{\star}$ in a halo of mass $M_{h}$, $n_{\star}(M_{\star}|z_{\ell})$ is the distribution of stellar mass at redshift $z_{\ell}$, and where we have defined $n_{\star}(\log M_{\star})=[M_{\star}/\log(e)]n_{\star}(M_{\star})$. Also, in the above, $dn/dM_{h}$ is the Halo Mass Function (HMF) for which we use Tinker2008 in the paper, and $n_{g}$ is the average number density of galaxies. In addition, in Equation (A) above we can write
@@
$\langle N(M_{h})\rangle$ $=\int_{z_{\ell,1}}^{z_{\ell,2}}dz_{\ell}\int_{M_{\star,1}}^{M_{\star,2}}d\log M_{\star}\;$
$\quad\times\langle N(M_{\star}|M_{h})\rangle n_{\star}(\log M_{\star}|z_{\ell})$
$=\int_{z_{\ell,1}}^{z_{\ell,2}}dz_{\ell}\int_{M_{\star,1}}^{M_{\star,2}}d\log M_{\star}\;$
$\quad\times\langle N(M_{\star}|M_{h})\rangle\frac{p(\log M_{\star},z_{\ell})}{n_{\ell}(z_{\ell})}\;,$ (51)
@@
$\quad\times\exp\left[-\left(\frac{\log M_{\star}-\log f_{\rm SHMR}(M_{h})}{\sigma_{\log M_{\star}}}\right)^{2}\right]\;.$ (52)
@@ Appendix B Validating the stellar mass catalog
We run two validation tests for our GalPro-generated stellar mass catalog based on Gneiting2007, which we summarize below. In order to perform these tests and create the plots we present in this section, we performs a $80/20$ split to create the training and validation sets, respectively, from the matched DES Y3 Deep Field and COSMOS2020 catalogs (see Section 4.1.1). GalPro is trained with the training set and the tests below are done with the validation set. See also Mucesh2020 for more details regarding the validation tests. More specifically, we test for probabilistic calibration, using the Probability Integral Transform (PIT), and for marginal calibration, by comparing the average predictive CDF and the true empirical CDF. We present the results, while briefly describing the process, for both tests below.
@@
    The results of the above tests on our model are shown in Figure 15. The PIT distributions follow closely the uniform. In addition, three statistics listed in the caption of the plot are given to quantitatively assess uniformity: the Kullback–Leibler (Kullback1951OnIA, KL) divergence, the Kolmogorov–Smirnov (Kolmogorov1992OnTE, KS) test, and the Cramér-von Mises (CramerVonMises1928, CvM) test. For each of these tests, values close to zero would indicate that our PIT distribution closely matches the Uniform distribution. We conclude that our model passes this test.
@@
    Although the PIT test is a necessary condition for the marginal PDF to be valid, it is usually not enough to make concrete statements about how well a PDF is calibrated. In fact, it has been shown (Hamill2001) that biased distributions can lead to uniformly distributed PIT’s. This test is thus not sufficient and we also perform the additional test below.
@@
    where we have defined the predictive CDF per galaxy as $F_{i}$ for all $n$ test galaxies i=1,2,,ni=1,2,...,n, and where
@@
Figure 15: Redshift (upper panel) and stellar-mass (lower panel) PIT distribution plots (number of objects NN versus PITPIT values) for the ML model trained to produce the DES Y3 stellar-mass sample. The quantile-quantile ($Q_{\rm theory}$$Q_{\rm data}$) plots are also shown in each plot to highlight deviations from uniformity. The black-dashed and solid blue lines represent the quantiles of the uniform and PIT distributions, respectively. The percentage of catastrophic outliers along with the values of the Kullback–Leibler (KL) divergence, Kolmogorov–Smirnov (KS) test, and Cramér-von Mises (CvM) metrics are also stated to quantify uniformity of the PIT distributions. Values close to zero for each of these statistics would indicate uniformity. We define a catastrophic outlier to be any galaxy with a redshift or stellar mass completely outside the support of its marginal PDF.
@@ Appendix C Computing the LSS weights
To prevent the network from overfitting, it is constructed with k-fold cross-validation, which works in the following way: The NSIDE = $4096$ maps are re-binned into a coarser grid of NSIDE = $16$ (with a resolution of about 4 degrees). We then randomly divide these cells into $k$ equal-area groups. To derive the weights for a given fold $k$, we train the neural network on the other folds, using fold $k$ as a validation sample (the training halts when the training metric no longer improves on the validation set). This cross-validation scheme will only work to prevent overfitting on scales below the resolution defined by NSIDE split, in this case around 4 degrees. As also mentioned in HighzSample, testing the corrected and uncorrected galaxy clustering on log-normal mock catalogs demonstrated no overfitting from this method at scales below 1 degree, and an impact of around 5% overfitting at scales of 2 degrees. We thus keep the galaxy clustering analysis in this work to angular scales below 1 degree to be conservative.
@@ Appendix D Validation of covariance matrix
As discussed in Section 5.2, our main analysis uses a Jackknife covariance, but we also calculate a theoretical covariance matrix via the CosmoCov (Krause2017_cosmolike; fang20_cosmocov2) package. Figure 17 shows the comparison between the Jackknife and CosmoCov error bars (the square root of the diagonal of the covariance matrix) in the case of lens-source bin combination $(\ell,m,s)=(2,1,4)$, for $\gamma_{t}$ (upper panel) and $w$ (lower panel); we use that bin for demonstration here but we see similar trends in all bins. In the left plot we also show the line $\propto 1/\theta$ which shows the expected behavior of the covariance if it were shape noise dominated. As we see in that plot, $\gamma_{t}$ is dominated by shape noise on small scales. This also implies that the cross-covariance between $\gamma_{t}$ and $w$ in a single bin and across different redshift and stellar-mass bins is weak. Therefore, we choose to ignore the cross-covariance in this work.
@@
The agreement between Jackknife and theory is not perfect. However, this is not very surprising given the differences in the HOD theory used by CosmoCov compared to this study, as the one implemented in the former is based on zheng2005; zehavi2011. Nevertheless, the differences we see in the covariance are not expected to have a significant effect on our data fits and we proceed to use the Jackknife covariace matrix in Section 6 while performing the likelihood analysis.